OpenStudy (anonymous):

If f(x) = x + 2, find the value of x if f(x) = 12.

OpenStudy (anonymous):

oh ok so all you do is fill in 12 for x do you get it?

jimthompson5910 (jim_thompson5910):

lizzyend515, that would be true if it said "x = 12" but it's saying "f(x) = 12"

jimthompson5910 (jim_thompson5910):

f(x) = x + 2 12 = x + 2 ... replace f(x) with 12 (since f(x) = 12) now solve for x

OpenStudy (anonymous):

then add 2

jimthompson5910 (jim_thompson5910):

you have +2 on the right side you have to undo it to isolate x what undoes +2?

OpenStudy (anonymous):

-

jimthompson5910 (jim_thompson5910):

yes, subtract 2 from both sides

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

good, x = 10

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome